OpenStudy (anonymous):

three times as many robins as cardinals visited a bird feeder. if a total of 20 robins and cardinals visited the feeder, how many were robins?

OpenStudy (texaschic101):

3c = r c + r = 20 sub the 3c in for r in the 2nd equation c + (3c) = 20 -- now combine like terms 4c = 20 -- now divide by 4 to get c by itself 4c/4 = 20/4 c = 5 now sub 5 in for c in either of the equations 3c = r 3(5) = r 15 = r check... c + r = 20 5 + 15 = 20 20 = 20 (correct) There were 15 robins and 5 cardinals.

OpenStudy (texaschic101):

Do you have any questions....if so....please ask and I can explain what I did.

OpenStudy (espex):

@texaschick101 The idea on OS is to help people learn not to simply do their problems for them. Give a man a fish and you feed him for a day, teach a man to fish and you feed him for life.